# How to Bridge the Sim-to-Real Gap in Digital Twin-Aided Telecommunication Networks

### v1

@@ I Introduction
A promising approach to address data scarcity in telecommunications is the use of digital twins (DTs)—high-fidelity simulators or emulators of network environments that can efficiently generate synthetic data tailored to specific deployments. By training AI models on a combination of real and synthetic data, it becomes possible to improve generalization performance compared to training solely on the limited real-world data typically available for a given deployment. Examples of DTs include NVIDIA’s Sionna library, which offers an efficient simulator of the propagation environment via ray tracing [1], with recent extensions also incorporating networking aspects [2]. DTs may also encompass real-world elements such as software-defined radios [3].
@@
How can we bridge the sim-to-real gap to ensure that AI models trained with DTs remain reliable in real deployments? As illustrated in Fig. 2(a), this article explores two complementary strategies:
@@
    Modeling the sim-to-real gap on the training objective (Sec. IV): Synthetic data produced by a DT are typically used to train AI models, e.g., the beamforming mapping in Fig. 1. In these settings, training is often done by minimizing a training loss evaluated using both real and synthetic data. In light of this, a complementary approach to devising reliable AI training strategies in the presence of DTs is to mitigate the sim-to-real gap directly at the level of the training loss. These techniques seek to reduce the DT bias under sim‑to‑real gaps by incorporating available real‑world observations. A prominent approach to achieve this is prediction-powered inference (PPI) [4] (see Fig. 2(c)).
@@ II Calibrating Digital Twins
Emerging DTs of telecommunication networks often include models of the wireless propagation environment using ray tracing [1]. In a ray tracing simulator, the environment is defined by a three-dimensional geometry along with a vector of material parameters describing the reflection, refraction, and diffraction characteristics of surfaces in the scene. The simulator models the propagation of radio waves by tracing the paths connecting transmitters to receivers, keeping track of the interactions of each path with the objects in the environment. The multi-path parameters returned by the simulator can then be leveraged to produce a simulated channel response or to estimate the received power at given locations (see Fig. 1).
In practice, however, a sim-to-real gap is inevitable: the specified geometry of the scene might miss objects or report inaccurate placements or dimensions, while material parameters may be unknown or wrongly specified. These mismatches lead to biases in the synthetic data generated by the DT. Therefore, it is important to reduce the sim-to-real gap caused by model misspecification by leveraging available real-world data to calibrate the DT.
@@ II-A Material and Parameter Calibration via Differentiable Simulation
One way to calibrate a DT is to learn the unknown material parameters by comparing simulation outputs to real measurements. To elaborate, we denote as $\rho$ the material parameters for the objects present in the scene. These include permittivities, conductivities, and scattering patterns. In practice, it is also possible to define an implicit neural representation (INR) that associates each position in the scene with a vector $\rho$ [1].
To support calibration, the network operator deploys a few test transmitters and receivers in the real network, and records the channel measurements $H$. Given a vector of model parameters $\rho$, one can then run the ray tracing simulator to yield a prediction $\hat{H}(\rho)$ of the channel measurements. Using this information, calibration can be formulated as an inverse problem: find the parameters $\hat{\rho}$ that cause the simulated measurement $\hat{H}(\hat{\rho})$ to best match the real observations $H$. This criterion is averaged over the available data set in order to obtain statistically sound estimates of the parameters $\rho$.
Recent research has shown that this inverse problem can be efficiently addressed by leveraging differentiable ray tracing software, and applying gradient-based optimization techniques to adjust the material parameters [1]. This approach treats the DT as a trainable model mapping material parameters to multi-path parameters, making it possible to differentiate through the ray tracing engine. This, in turn, supports the use of gradient descent to infer material properties $\rho$ that minimize the error between simulated and measured channel responses.
@@ II-B Phase Error-Aware Calibration
Calibrating material parameters is useful, but it cannot fully resolve biases caused by a misspecified geometry of the objects in the scene. For instance, as seen in Fig. 2(a), minor errors in building shapes or additional objects not present in the DT can cause significant phase errors in the simulated propagation paths. In fact, even a small displacement of an object—of the order of a fraction of the wavelength—can cause the phase of a reflected path to shift, leading to constructive or destructive interference patterns that do not align with reality [5]. Owing to these unaccounted phase errors, traditional calibration methods that only tweak material values $\rho$ may fail to produce an accurate simulator.
To tackle this issue, a promising approach is to devise calibration schemes that are aware of the presence of phase errors caused by a sim-to-real gap in the geometry of the environment. The core idea behind phase error-aware calibration, as proposed in [5], is to iteratively estimate and correct phase errors in the simulated channel. In this method, the DT’s prediction is augmented with per-path phase error terms that account for unmodeled geometry nuances. An expectation-maximization (EM) algorithm is then used: in the E-step, the algorithm infers the likely phase errors for each propagation path by comparing the simulation with real measurements; while the M-step updates the simulator parameters $\rho$ to better fit the observed data, effectively absorbing the phase errors into the model.
Fig. 3 shows the performance of different calibration techniques for a canyon scenario with line-of-sight (LoS) blockage under geometry mismatch at the DT [5]. The calibration procedures are conducted under multiple levels of available bandwidth for channel measurements, and the results quantify the relative error between the measured power and the power predicted by the calibrated DT. We compare the described phase error-aware approach with the following alternative methods: (i) phase error-oblivious calibration, which disregards phase errors due to geometry mismatch when matching simulated and measured channel responses; and (ii) uniform phase error calibration, which assumes the phases of the predicted paths to be uniformly distributed when matching delay-angle projections of the predicted and measured channel powers during calibration. All methods tune the material parameters $\rho$ via gradient-based optimization [1] (see Sec. II-A).
@@
Figure 3: Relative power prediction errors of ray tracing-based DTs calibrated using phase error-oblivious (orange), uniform phase error (green), and phase error-aware (blue) calibration methods, where calibration is conducted using channel measurements at different bandwidths (figure adapted from [5]).
@@ III Modeling the Sim-to-Real Gap on the Environment
Figure 4: Training a multiple access protocol using synthetic data from a digital twin (DT) modeling the access channel: (left) a Bayesian formulation of DT learning assigns different configurations (interference levels) of the DT to a probability distribution that depends on prior knowledge and available data. In contrast, a frequentist approach would only select the highest-probability model. (center) While the frequentist DT trains the policy inside the single selected model, the Bayesian DT generates data from multiple models, weighting the importance of multiple plausible DT models during policy optimization. (right) When calibration data is in limited supply, explicitly modeling the sim-to-real gap via uncertainty-aware Bayesian methods can significantly improve the performance of models trained using synthetic data from the DT (adapted from [6]).
@@
As an exemplifying use case, following [6], we now focus on a system in which multiple devices coordinate to access a shared wireless medium using a protocol learned via multi-agent reinforcement learning (MARL). The MARL protocol is trained using synthetic data from a DT. If we were to use a single DT model to simulate the environment, as done in a conventional frequentist implementation, the resulting policy might perform well under that specific model, but it could fail to generalize to the real-world system due to the real-to-sim gap. In contrast, the Bayesian approach samples multiple plausible environments from the posterior distribution, training the MARL policy across a number of different DT models. This leads to a policy that is robust under the sim-to-real gap modeled at the DT, achieving better generalization to the true environment.
@@
Fig. 4 elaborates on the operation and on the advantages of the Bayesian methodology for the given multiple access setting. The left panel illustrates a posterior distribution over different DT models that assume varying levels of interference, which result in distinct likelihoods to the number of received packets given the number of concurrently transmitted packets. The central panel contrasts the training processes of different methodologies. The conventional frequentist learning approach relies on a single maximum a posteriori model, while the Bayesian learning framework integrates across a range of likely models, weighting each model according to its posterior probability. As shown in the right panel, this yields superior performance for the policy trained using synthetic data, especially in data-scarce regimes where the model uncertainty about the environment dynamics is high.
@@
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [7, Ch. 12]. It should be emphasized that, while these approaches improve robustness and reliability, they do come at the cost of increased computational complexity: sampling multiple environments, training across a population of models, and maintaining distributions over parameters may introduce significant overhead. Nonetheless, in many critical applications where reliability and safety are paramount, this additional cost often offers a worthwhile trade-off.
@@ IV Modeling the Sim-to-Real Gap on the Training Objective
Another approach to address the residual sim-to-real gap is to leave the DT as is, but modify the training algorithm to compensate for the DT’s bias. A recently developed framework called PPI [4, 8] offers a powerful tool in this regard. PPI originates from the idea of doubly robust estimation in statistics, and it supports the use of abundant biased data from a simulator, together with a small amount of unbiased data from the real world, to produce unbiased learning outcomes.
@@ IV-A Semi-Supervised Learning via Prediction-Powered Inference
PPI supports the definition of a semi-supervised training methodology that leverages both real and synthetic data. To elaborate, suppose we want to train an AI model for a certain task using mainly synthetic data generated by the DT. Assume also that we also have a limited set of real data for the same task, but not enough to train a high-capacity model from scratch. In classical transfer learning, one might pre-train on synthetic data and then fine-tune on real data [7, 9].
In contrast, as illustrated in Fig. 2(c), PPI uses the real data not to train the model directly, but to estimate the bias of a training loss estimate based on synthetic data. It then adjusts the model training objective to remove this bias. Essentially, the real data provides a yardstick for how far off the synthetic training loss is from the real training loss. PPI is appealing for telecommunication AI problems because obtaining even a small validation set of real data is often feasible through field tests or pilot deployments, whereas gathering a comprehensive training set is not [10].
The corrected training objective produced by PPI is guaranteed to be unbiased with respect to the real data distribution. In this regard, a key advantage is that the real data are used very efficiently, as they are not wasted on brute-force training, but rather on pinpointing the bias. This approach reduces reliance on real data by leveraging less expensive synthetic samples, enabling models to perform well in practice even when only a small fraction, e.g. $1$-$10\%$, of the training data is real [10].
(a) Cross-PPI
$2\cdot 10^{-2}$$4\cdot 10^{-2}$$6\cdot 10^{-2}$$8\cdot 10^{-2}$0.10.10.10.13.53.53.53.544444.54.54.54.555555.55.55.55.5Proportion of real dataChannel capacity (bps/Hz)ERMP-ERMPPICross-PPI
(b) Cross-PPI on beam alignment via CKM
Figure 5: (a) Illustration of Cross-PPI [8, 10]: The real data is divided into $K$ folds, and $K$ calibration procedures are conducted, each using all real data except one fold, held out for bias estimation. The calibration bias is estimated and removed from the synthetic data loss (as shown in Fig. 2(c)), obtaining the Cross-PPI loss. (b) Channel capacity as a function of the proportion of real data for Cross-PPI, PPI [4], empirical risk minimization (ERM), and pseudo-empirical risk minimization (P-ERM). As detailed in [10], given real data combined with synthetic samples, a classification model is trained to map UE-location to the optimal downlink beam index. The performance is measured in terms of achievable downlink channel capacity.
That said, PPI in its basic form has two important limitations when it comes to applications to telecommunication systems:
$\bullet$ Data efficiency: The requirement to hold out a portion of real data exclusively for bias correction means that those data cannot be used to calibrate the DT. If real data are extremely scarce, this may be a significant drawback.
$\bullet$ Context independence: The bias correction in vanilla PPI is typically a single global adjustment. In other words, it assumes the simulator’s bias is the same in all situations. In telecommunication contexts, the accuracy of a DT may vary with the scenario. For example, the DT might be very accurate for urban microcell cases, but less so for rural macrocells, or accurate for certain ranges of signal-to-noise ratio but not others [11]. A single bias term would struggle to correct all contexts equally.
@@ IV-B Cross-Validation-Based Prediction-Powered Inference
Cross-PPI improves data efficiency by effectively recycling real data during calibration while still using it for bias estimation. As depicted in Fig. 5(a), this is achieved through a cross-validation-style procedure in which the real data is partitioned into $K$ folds, and synthetic data is calibrated using $K$ calibration procedures, where each procedure uses all real data except one fold, reserved for bias correction [8, 10]. In such a way, all real data is used for both calibration and bias correction.
To illustrate the advantages of Cross-PPI, we consider the problem of beam alignment via channel knowledge maps (CKMs) in mmWave massive MIMO systems [12]. A CKM is a site‑specific database of channel characteristics associated with the locations of both the transmitter and the receiver [12]. Recently, this concept has been leveraged to mitigate training overhead, offering a promising new approach for channel estimation and beam alignment [12].
The objective is to train a classification model that maps the UE location to the optimal downlink beam selected from a predefined finite codebook. Fig. 5(b) shows the performance of the trained model in terms of achievable downlink channel capacity as a function of the fraction of real samples. Three baselines are considered: empirical risk minimization (ERM), which ignores the synthetic data; pseudo-empirical risk minimization (P-ERM), which treats the synthetic data as ground‑truth samples; and PPI [4]. The results indicate that Cross‑PPI achieves substantial improvements compared to all other methods. Notably, P-ERM does not yield any clear benefits over ERM, which relies exclusively on the real data. Although PPI shows some improvements over the baselines, its performance remains lower than that of Cross‑PPI.
@@
Figure 6: (a) Heatmap for the ground-truth optimal azimuth angles required to serve a user in the given position from the BS placed at the black cross mark. The layouts of the buildings are depicted as empty white regions. (b) Optimal azimuth angles as prediced by a DT. (c)-(e) Heatmap for the pointing error obtained by a model trained using real and synthetic data via (c) empirical risk minimization (ERM), (d) PPI [13], and (e) context-aware PPI (CPPI) [11].

### v2

@@ I Introduction
A promising approach to address data scarcity in telecommunications is the use of digital twins (DTs)—high-fidelity simulators or emulators of network environments that can efficiently generate synthetic data tailored to specific deployments. By training AI models on a combination of real and synthetic data, it becomes possible to improve generalization performance compared to training solely on the limited real-world data typically available for a given deployment. Examples of DTs include NVIDIA’s Sionna library, which offers an efficient simulator of the propagation environment via ray tracing [hoydis2024learning], with recent extensions also incorporating networking aspects [pegurri2024toward]. DTs may also encompass real-world elements such as software-defined radios [polese2024colosseum].
@@
How can we bridge the sim-to-real gap to ensure that AI models trained with DTs remain reliable in real deployments? As illustrated in Fig. 2, this article explores two complementary strategies:
@@
    Modeling the sim-to-real gap on the training objective (Sec. IV): Synthetic data produced by a DT are typically used to train AI models, e.g., the beamforming mapping in Fig. 1. In these settings, training is often done by minimizing a training loss evaluated using both real and synthetic data. In light of this, a complementary approach to devising reliable AI training strategies in the presence of DTs is to mitigate the sim-to-real gap directly at the level of the training loss. These techniques seek to reduce the DT bias under sim‑to‑real gaps by incorporating available real‑world observations. A prominent approach to achieve this is prediction-powered inference (PPI) [angelopoulos2023prediction] (see Fig. 2(c)).
As a final remark, we emphasize that this paper focuses on leveraging DTs for data augmentation in AI model training. While this represents an important use case, as highlighted by O-RAN [oran2024dtran, li2025generative], DTs can also enable real-time operational tasks, including, for instance, channel and link quality prediction [DTrev]. Such real-time applications require continuous synchronization mechanisms to manage the real-to-sim gap at fine temporal granularities [hashash2022edge]. In contrast, our work addresses DT calibration at a non-real-time scale, focusing on training AI models using DTs that capture long-term environmental statistics and operating conditions.
@@ II Calibrating Digital Twins
Emerging physics-based DTs of telecommunication networks often include models of the wireless propagation environment using ray tracing [hoydis2024learning]. In a ray tracing simulator, the environment is defined by a three-dimensional geometry along with a vector of material parameters describing the reflection, refraction, and diffraction characteristics of surfaces in the scene. The simulator models the propagation of radio waves by tracing the paths connecting transmitters to receivers, keeping track of the interactions of each path with the objects in the environment. The multi-path parameters returned by the simulator can then be leveraged to produce a simulated channel response or to estimate the received power at given locations (see Fig. 1).
In practice, however, a sim-to-real gap is inevitable: the specified geometry of the scene might miss objects or report inaccurate placements or dimensions, while material parameters may be unknown or wrongly specified. These mismatches lead to biases in the synthetic data generated by the DT. Therefore, as discussed in this section, it is important to reduce the sim-to-real gap caused by model misspecification by leveraging available real-world data to calibrate the DT.
While this section focuses on the calibration of physics-based models, data-driven models, such as dynamic systems, can also benefit from periodic calibration based on current data. This can be done using standard continual learning tools, which are not covered here [hashash2022edge].
@@ II-A Material and Parameter Calibration via Differentiable Simulation
One way to calibrate a DT is to learn the unknown material parameters by comparing simulation outputs to real measurements. To elaborate, we denote as $\rho$ the material parameters for the objects present in the scene. These include permittivities, conductivities, and scattering patterns. In practice, it is also possible to define an implicit neural representation (INR) that associates each position in the scene with a vector $\rho$ [hoydis2024learning].
[…]
Recent research has shown that this inverse problem can be efficiently addressed by leveraging differentiable ray tracing software, and applying gradient-based optimization techniques to adjust the material parameters [hoydis2024learning]. This approach treats the DT as a trainable model mapping material parameters to multi-path parameters, making it possible to differentiate through the ray tracing engine. This, in turn, supports the use of gradient descent to infer material properties $\rho$ that minimize the error between simulated and measured channel responses.
@@ II-B Phase Error-Aware Calibration
Calibrating material parameters is useful, but it cannot fully resolve biases caused by a misspecified geometry of the objects in the scene. For instance, as seen in Fig. 2(a), minor errors in building shapes or additional objects not present in the DT can cause significant phase errors in the simulated propagation paths. In fact, even a small displacement of an object—of the order of a fraction of the wavelength—can cause the phase of a reflected path to shift, leading to constructive or destructive interference patterns that do not align with reality [ruah2024calibrating]. Owing to these unaccounted phase errors, traditional calibration methods that only tweak material values $\rho$ may fail to produce an accurate simulator.
To tackle this issue, a promising approach is to devise calibration schemes that are aware of the presence of phase errors caused by a sim-to-real gap in the geometry of the environment. The core idea behind phase error-aware calibration, as proposed in [ruah2024calibrating], is to iteratively estimate and correct phase errors in the simulated channel. In this method, the DT’s prediction is augmented with per-path phase error terms that account for unmodeled geometry nuances. An expectation-maximization (EM) algorithm is then used: in the E-step, the algorithm infers the likely phase errors for each propagation path by comparing the simulation with real measurements; while the M-step updates the simulator parameters $\rho$ to better fit the observed data, effectively absorbing the phase errors into the model.
Fig. 3 shows the performance of different calibration techniques for a canyon scenario with line-of-sight (LoS) blockage, resulting in two propagation paths connecting the transmitter to the receiver by reflecting on each of the canyon’s walls. In this setting, geometry mismatch at the DT is induced by a misplacement of one of the walls with respect to the true scene [ruah2024calibrating]. The calibration procedures are conducted under multiple levels of available bandwidth for channel measurements, and the results quantify the relative error between the measured power and the power predicted by the calibrated DT. We compare the described phase error-aware approach with the following alternative methods: (i) phase error-oblivious calibration, which disregards phase errors due to geometry mismatch when matching simulated and measured channel responses; and (ii) uniform phase error calibration, which assumes the phases of the predicted paths to be uniformly distributed when matching delay-angle projections of the predicted and measured channel powers during calibration. All methods tune the material parameters $\rho$ via gradient-based optimization [hoydis2024learning] (see Sec. II-A).
@@
Figure 3: Relative power prediction errors of ray tracing-based DTs calibrated using phase error-oblivious (orange), uniform phase error (green), and phase error-aware (blue) calibration methods, where calibration is conducted using channel measurements at different bandwidths (figure adapted from [ruah2024calibrating]).
Figure 4: Training a multiple access protocol using synthetic data from a digital twin (DT) modeling the access channel: (left) a Bayesian formulation of DT learning assigns different configurations (interference levels) of the DT to a probability distribution that depends on prior knowledge and available data. In contrast, a frequentist approach would only select the highest-probability model. (center) While the frequentist DT trains the policy inside the single selected model, the Bayesian DT generates data from multiple models, weighting the importance of multiple plausible DT models during policy optimization. (right) When calibration data is in limited supply, explicitly modeling the sim-to-real gap via uncertainty-aware Bayesian methods can significantly improve the performance of models trained using synthetic data from the DT (adapted from [ruah2023bayesian]).
@@ III Modeling the Sim-to-Real Gap on the Environment
As an exemplifying use case, following [ruah2023bayesian], we now focus on a system in which multiple devices coordinate to access a shared wireless medium using a protocol learned via multi-agent reinforcement learning (MARL). The MARL protocol is trained using synthetic data from a DT. If we were to use a single DT model to simulate the environment, as done in a conventional frequentist implementation, the resulting policy might perform well under that specific model, but it could fail to generalize to the real-world system due to the real-to-sim gap. In contrast, the Bayesian approach samples multiple plausible environments from the posterior distribution, training the MARL policy across a number of different DT models. This leads to a policy that is robust under the sim-to-real gap modeled at the DT, achieving better generalization to the true environment.
@@
Fig. 4 elaborates on the operation and on the advantages of the Bayesian methodology for a simulated multiple access setting composed of four devices that communicate with a single base station (BS) under an unknown shared channel. The shared channel attributes lower probabilities of success to the delivery of packets when multiple devices transmit within the same time slot [ruah2023bayesian]. Accordingly, the left panel of Fig. 4 illustrates a posterior distribution over different DT models that assume varying levels of interference, resulting in distinct likelihoods to the number of received packets given the number of concurrently transmitted packets. The central panel contrasts the training processes of different methodologies. The conventional frequentist learning approach relies on a single maximum a posteriori model, while the Bayesian learning framework integrates across a range of likely models, weighting each model according to its posterior probability. As shown in the right panel, this yields superior performance for the policy trained using synthetic data, especially in data-scarce regimes where the model uncertainty about the environment dynamics is high.
@@
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [simeone2022machine, Ch. 12]. It should be emphasized that, while these approaches improve robustness and reliability, they do come at the cost of increased computational complexity: sampling multiple environments, training across a population of models, and maintaining distributions over parameters may introduce significant overhead. Nonetheless, in many critical applications where reliability and safety are paramount, this additional cost often offers a worthwhile trade-off.
(a) Cross-PPI
$2\cdot 10^{-2}$$4\cdot 10^{-2}$$6\cdot 10^{-2}$$8\cdot 10^{-2}$0.10.13.53.5444.54.5555.55.5Proportion of real dataChannel capacity (bps/Hz)ERMP-ERMPPICross-PPI
(b) Beam alignment via CKM
Figure 5: (a) Illustration of Cross-PPI [sifaou2024semi]: The real data is divided into $K$ folds, and $K$ calibration procedures are conducted, each using all real data except one fold, held out for bias estimation. The calibration bias is estimated and removed from the synthetic data loss (as shown in Fig. 2(c)), obtaining the Cross-PPI loss. (b) Channel capacity as a function of the proportion of real data for Cross-PPI, PPI [angelopoulos2023prediction], empirical risk minimization (ERM), and pseudo-empirical risk minimization (P-ERM). As detailed in [sifaou2024semi], given real data combined with synthetic samples, a classification model is trained to map UE-location to the optimal downlink beam index. The performance is measured in terms of achievable downlink channel capacity.
@@ IV Modeling the Sim-to-Real Gap on the Training Objective
Another approach to address the residual sim-to-real gap is to leave the DT as is, but modify the training algorithm to compensate for the DT’s bias. In this regard, a recently developed framework called PPI [angelopoulos2023prediction] offers a powerful tool for training AI models when the amount of real world data is too small to reach the desired level of accuracy. PPI originates from the idea of doubly robust estimation in statistics, and it supports the use of abundant biased data from a simulator, together with a small amount of unbiased data from the real world, to produce unbiased learning outcomes.
@@ IV-A Semi-Supervised Learning via Prediction-Powered Inference
PPI supports the definition of a semi-supervised training methodology that leverages both real and synthetic data. To elaborate, suppose we want to train an AI model for a certain task using mainly synthetic data generated by the DT. Assume also that we also have a limited set of real data for the same task, but not enough to train a high-capacity model from scratch. In classical transfer learning, one might pre-train on synthetic data and then fine-tune on real data [simeone2022machine].
In contrast, as illustrated in Fig. 2(c), PPI uses the real data not to train the model directly, but to estimate the bias of a training loss estimate based on synthetic data. It then adjusts the model training objective to remove this bias. Essentially, the real data provides a yardstick for how far off the synthetic training loss is from the real training loss. PPI is appealing for telecommunication AI problems because obtaining even a small validation set of real data is often feasible through field tests or pilot deployments, whereas gathering a comprehensive training set is not [sifaou2024semi].
The PPI-corrected training objective is unbiased with respect to the real data distribution and uses real data efficiently: instead of brute-force training, real samples are used to estimate simulator bias, while cheaper synthetic data provide most of the training signal. This enables good performance even when only a small fraction, e.g. 11-$10\%$, of the training data is real [sifaou2024semi]. That said, PPI in its basic form has two key limitations for telecommunication systems:
$\bullet$ Data efficiency: A subset of real data must be reserved solely for bias correction and cannot be used to calibrate the DT, which is problematic when real data are extremely scarce.
$\bullet$ Context independence: Bias is typically corrected by a single global term, implicitly assuming the simulator error is the same in all contexts. In practice, DT accuracy may vary by scenario, e.g., urban microcells vs. rural macrocells or across different signal-to-noise ratio ranges [ruah2025context], so a single bias term cannot adequately correct all contexts.
@@ IV-B Cross-Validation-Based Prediction-Powered Inference
Cross-PPI improves data efficiency by effectively recycling real data during calibration while still using it for bias estimation. As depicted in Fig. 5(a), this is achieved through a cross-validation-style procedure in which the real data is partitioned into $K$ folds, and synthetic data is calibrated using $K$ calibration procedures, where each procedure uses all real data except one fold, reserved for bias correction [sifaou2024semi]. In such a way, all real data is used for both calibration and bias correction.
To illustrate the advantages of Cross-PPI, we consider the problem of beam alignment via channel knowledge maps (CKMs) in mmWave massive MIMO systems [zeng2021toward]. A CKM is a site‑specific database of channel characteristics associated with the locations of both the transmitter and the receiver [zeng2021toward]. Recently, this concept has been leveraged to mitigate training overhead, offering a promising new approach for channel estimation and beam alignment [zeng2021toward].
The objective is to train a classification model that maps the UE location to the optimal downlink beam selected from a predefined finite codebook. Fig. 5(b) shows the performance of the trained model in terms of achievable downlink channel capacity as a function of the fraction of real samples. Three baselines are considered: empirical risk minimization (ERM), which ignores the synthetic data; pseudo-empirical risk minimization (P-ERM), which treats the synthetic data as ground‑truth samples; and PPI [angelopoulos2023prediction]. The results indicate that Cross‑PPI achieves substantial improvements compared to all other methods. Notably, P-ERM does not yield any clear benefits over ERM, which relies exclusively on the real data. Although PPI shows some improvements over the baselines, its performance remains lower than that of Cross‑PPI.
@@
Figure 6: (a) Heatmap for the ground-truth optimal azimuth angles required to serve a user in the given position from the BS placed at the black cross mark. The layouts of the buildings are depicted as empty white regions. (b) Optimal azimuth angles as predicted by a DT. (c)-(e) Heatmap for the pointing error obtained by a model trained using real and synthetic data via (c) empirical risk minimization (ERM), (d) PPI [angelopoulos2023prediction], and (e) context-aware PPI (CPPI) [ruah2025context].